OpenStudy (anonymous):

Solve the given equation for n. 12(n+5) = 3 (20+4n) Select the correct choice below and, if necessary, fill in the answer box to complete your choice. A. The solution is n= B. The solution is all real numbers C. There is no solution

Nnesha (nnesha):

distribute first bracket by 12 and second by 3 let me know what you get after that

Nnesha (nnesha):

n times 12=?? 5 times 12=??? 20 times 3=??? 4n time 3 = ???

OpenStudy (xxdragoonxx):

12(n+5)=3(20+4n) 4(n+5)=20+4n = 4n+20=20+4n = 20=20+4n-4n = 20+4n-4n=20

OpenStudy (anonymous):

Then what do I do? @xxDragoonxx

OpenStudy (xxdragoonxx):

Many solutions.

OpenStudy (anonymous):

n=40 correct? @xxDragoonxx

OpenStudy (xxdragoonxx):

20.

OpenStudy (anonymous):

It said that was wrong -.-

OpenStudy (anonymous):

It said 20 was wrong too @xxDragoonxx

Nnesha (nnesha):

because you have to distribute then you can see that n should cancel out

Nnesha (nnesha):

 now